Mode Cleaner doesn't want to stay locked. Seismic is coming down from an earthquake ~20min ago.
We're in the process of measuring IPPOS, so this is obnoxious.
EDIT: Followed by a 6.2 and a 7.1 at 07:06UTC and 07:15UTC in the same area.
We're following the tried and true tradition of going home when there's an earthquake big enough that the MC won't stay locked.
Several optics have rung up, PRM is the only one which has tripped so far, because the side sensor has the extra gain, but the watchdog threshold is set for the face OSEMs.